# Proc2Pddl: Open-Domain Planning Representations from Texts

### v1

@@ Abstract
Planning in a text-based environment continues to be a major challenge for AI systems. Recent approaches have used language models to predict a planning domain definition (e.g., PDDL) but have only been evaluated in closed-domain simulated environments. To address this, we present Proc2Pddl, the first dataset containing open-domain procedural texts paired with expert-annotated PDDL representations. Using this dataset, we evaluate state-of-the-art models on defining the preconditions and effects of actions. We show that Proc2Pddl is highly challenging, with GPT-3.5’s success rate close to 0% and GPT-4’s around 35%. Our analysis shows both syntactic and semantic errors, indicating LMs’ deficiency in both generating domain-specific prgorams and reasoning about events. We hope this analysis and dataset helps future progress towards integrating the best of LMs and formal planning. [^1] Our resources can be found at https://github.com/zharry29/nl-to-pddl-action.
@@
[…]
@@ 1 Introduction
Planning is the task of finding a sequence of actions to achieve a goal in a given environment Fikes and Nilsson (1971); LaValle (2006). In real life, the environment is often described with natural language texts. To enable text-based, automated planning, recent work has used language models (LMs) to generate plans Valmeekam et al. (2023a); Stein and Koller (2023). However, this approach is found to fall short with regard to both performance and interpretability Valmeekam et al. (2023c, b). Alternatively, another recent line of worked has instead used LMs to translate the natural language description of environments to planning domain definition language (PDDL) Ghallab et al. (1998). This symbolic representation can then be solved by a planner in a plan Collins et al. (2022); Lyu et al. (2023); Liu et al. (2023); Xie et al. (2023); Wong et al. (2023). Despite of success of such a neurosymbolic method, all the above work has only been evaluated in closed-domains simulated environments such as a household (e.g., ALFRED Shridhar et al. (2020)) or discrete object placement (e.g., BlocksWorld Valmeekam et al. (2024)) (see Table 1).
@@
Figure 2: Our formulation of the $\mathbb{DF}$-action prediction task. Given the domain file header that specifies the ontology, a model predicts action definitions including parameters, preconditions, and effects based on textual descriptions. During evaluation, the predicted $\mathbb{DF}$ is both compared to a reference and used to solve corresponding $\mathbb{PF}$s.
To enable open-domain, text-based planning, we propose Proc2Pddl, a dataset to evaluate models’ ability to generate PDDL given procedural texts. Proc2Pddl consists of 27 pairs of open-domain procedures and PDDL representations. Each PDDL representation include a domain file $\mathbb{DF}$ that models the types, predicates, and actions, and a problem file $\mathbb{PF}$ that models the entities, initial states, and goal states (see Figure 1). Because Proc2Pddl is not bound to any simulation, the PDDL representations are manually annotated by experts trained on this task to ensure validity, resulting in 27 domain files and 81 problem files.
Using this dataset, we study the task of action modeling Lindsay et al. (2017) formulated as follows. The input is some natural language texts and the header of a $\mathbb{DF}$ (i.e., types, predicates, and names of actions). The output is the definitions of actions in the $\mathbb{DF}$ (i.e., parameters, preconditions, and effects) (see Figure 2). During evaluation, the predicted $\mathbb{DF}$ and a ground-truth $\mathbb{PF}$ are provided to a PDDL solver that may or may not successfully produce a plan. We then evaluate the existence and correctness of such a plan. In this formulation, our assumption of the $\mathbb{DF}$ header is necessary to ensure the consistency of semantics between the $\mathbb{DF}$ and the $\mathbb{PF}$ for evaluation. It is also empirically motivated; for example, a kitchen robot may have access to the types like ‘ingredients’ and predicates like ‘diced’ via some information extraction system given descriptive texts, but it may still need to predict, for “swinging a knife”, the precondition that it is only safe to do so to the ‘ingredients’ and the effect that they will become ‘diced’.
Through our experiment, we show that the task of action modeling in Proc2Pddl is highly challenging to state-of-the-art LMs, where GPT-3.5 almost fails completely and GPT-4 can only generate exactly matching $\mathbb{DF}$s 16% of the time and solvable $\mathbb{DF}$s 33% of the time. By devising a summarize-extract-translate pipeline that prompt LMs to modularly generate PDDL, we improve both metrics by 2-3%. In our analysis, the syntactic errors indicate LMs’ weakness in generating low-resource and domain-specific programming languages Cassano et al. (2023) like PDDL, while the semantic errors suggest LMs’ inaccuracies to reason about actions and environments.
@@
2 Problem Formulation
The problem of predicting a planning domain definition in a text-based environment can be seen as translating natural language texts to PDDL, which consists of a domain file $\mathbb{DF}$ and one or more problem files $\mathbb{PF}$s.
A $\mathbb{DF}$ defines the following elements:
  • a header $H$, which consists of
    • types of entities (e.g., object, location, player)
    • predicates (e.g., if object is at a location)
    • names of possible actions (e.g., boil water)
  • definitions of actions $A$, which consist of
@@ A $\mathbb{DF}$ defines the following elements:
A $\mathbb{PF}$ defines the following elements:
  • entities and their type (e.g., rainwater is water)
@@ A $\mathbb{PF}$ defines the following elements:
Traditionally, the task of text-based action modeling involves predicting $\mathbb{DF}$ based on text $\mathbb{T}$, where a successfully generated $\mathbb{DF}$ can thus solve provided $\mathbb{PF}$s defined accordingly. In this paper we tackle a simplified formulation where the types, predicates, and names of actions ($H$, domain header) are provided; a model needs to predict the parameters, preconditions, and effects of each action ($A$, action definitions). As the information specified by $H$ is guaranteed to be consistent with that of the $\mathbb{PF}$s, the evaluation is well-defined.
3 Experiments
We provide some first attempts towards our formulation of action modeling in open-domain texts.
Dataset   We introduce the Proc2Pddl dataset of 27 different $\mathbb{T}$-$\mathbb{DF}$-$\mathbb{PF}$s tuples, drawing procedural texts from wikiHow articles of various topics (see Appendix A). A class of graduate students in a U.S. university with prior knowledge of PDDL are each given a wikiHow article and annotate a $\mathbb{DF}$ and multiple corresponding $\mathbb{PF}$s from the article, each with a gold plan to solve it. On average, there are 13.33 defined actions in a $\mathbb{DF}$ and 8.07 instantiated actions in a gold plan. During prediction, we treat $\mathbb{T}$ as an annotated one-line summary of all annotated actions, though results using more variations of $\mathbb{T}$ can be found in Appendix C. On average, it takes several hours to train each human annotator, and another several hours to produce a $\mathbb{T}$-$\mathbb{DF}$-$\mathbb{PF}$s tuple. We partition the 27 examples into a 5:6:16 train-development-test splits. In this work, the train split is unused as all our methods are zero-shot; only the development set is used for error analysis; the test set is strictly held out for evaluation. Some sample data of Proc2Pddl can be found in Appendix B.
Models   We focus on zero-shot prompting LMs as PDDL for any domain is extremely costly to annotate, leading to hardly any training data. To predict action definitions $A$ in $\mathbb{DF}$ based on the header $H$ and some $\mathbb{T}$, we prompt an LM in a zero-shot manner Brown et al. (2020) by describing the task and providing the input. Note that few-shot prompting is prohibitively costly due to the excessive length of a resulting $\mathbb{DF}$. We also incorporate the chain-of-thought (CoT) technique Wei et al. (2023) that explicitly prompts an LM to perform three essential sub-tasks:
  • Summarization: describe each action, including the expected preconditions and effects;
  • Extraction: list the involved entities and their states before and after;
@@ 3 Experiments
Model % action acc. $\mathbb{PF}$ solve exact plan
gpt-3.5 0.2 1.0 1.0
gpt-4 15.9 33.7 4.2
gpt-4 + CoT 18.1 35.8 6.3
gold 100.0 100.0 100.0
Table 2: LMs are greatly challenged by Proc2Pddl, where gpt-3.5 has close to zero performance, gpt-4 has non-trivial performance, while gpt-4 with chain-of-thought is slightly better.
@@
gpt-4 + CoT 42.2 29.7 48.1
Table 3: LMs (both GPT-4 models) are worse at predicting preconditions than effects.
The verbatim prompt wordings can be found in Appendix D. We experiment with two large LMs, gpt-3.5-turbo-16k and gpt-4-32k dated June 2023 with max tokens of 8192819281928192222Due to the need for very long input and output, the choice of open-source models is limited. We are in progress of implementing Mixtral-8x7B., temperature of $0.5$, and default hyperparameters otherwise.
@@ 4 Evaluation and Analysis
Now that a model generates the parameters, preconditions, and effects for each action, we have a complete $\mathbb{DF}$. We evaluate it in two ways (Figure 2). Intrinsically, we semantically compare the predicted $A$ with the ground-truth provided by our Proc2Pddl and report an action-wide accuracy. Equivalence of two action definitions does not depend on the naming of variables nor on the order within conjunctions. Extrinsically, to measure actions’ coherence, we use a BFS-based PDDL solver [^3] https://github.com/pucrs-automated-planning/pddl-parser to attempt to solve ground-truth $\mathbb{PF}$s with the predicted $\mathbb{DF}$ and report a success rate. An unsolved $\mathbb{PF}$ is caused by (1.) no plan can be found, or (2.) the solver runs for more than 30 seconds, or (3.) the solver returns an error (usually a syntax error in generated PDDL).
The intrinsic and extrinsic results are shown in Table 2. gpt-3.5-turbo which achieves impressive performance on many tasks has a close-to-zero performance. In contrast, its predecessor gpt-4 performs significantly better with 18% action prediction accuracy and 36% solve rate of $\mathbb{PF}$s based on the predicted $\mathbb{DF}$. Still, the performance far worse than ideal, showing that even a simplified open-domain planning formulation is challenging to state-of-the-art LMs.
CoT is helpful overall since it explicitly spells out many implicit entities and state changes in the extraction stage which are critical to predicting preconditions and effects. In most situations, the model summarizes the action and extracts the entity states correctly, though sometimes missing a few implicit entities. However, CoT’s bottleneck lies in the translation stage, during which there are mainly three types of errors.
@@
To address these, we leave to future work to demonstrate and standardize the translation process by clearly describing when an entity change or stage change is and is not needed, while also encouraging the model to strictly match the given predicates. Finer-grained evaluation results are shown in Table 3 to tease out the performance regarding such component within an action. It is clear that the LM is worse at predicting preconditions than at predicting effects. This is understandable as procedural texts like wikiHow tend to be less explicit about predictions than about effects (e.g., from bake for 10 minutes it is obvious that the food will be baked, but it is unclear what state it had been in).

### v2

@@ Abstract
Planning in a text-based environment continues to be a significant challenge for AI systems. Recent approaches have utilized language models to predict planning domain definitions (e.g., PDDL) but have only been evaluated in closed-domain simulated environments. To address this, we present Proc2Pddl, the first dataset containing open-domain procedural texts paired with expert-annotated PDDL representations. Using this dataset, we evaluate the task of predicting domain actions (parameters, preconditions, and effects). We experiment with various large language models (LLMs) and prompting mechanisms, including a novel instruction inspired by the zone of proximal development (ZPD), which reconstructs the task as incremental basic skills. Our results demonstrate that Proc2Pddl is highly challenging for end-to-end LLMs, with GPT-3.5’s success rate close to 0% and GPT-4o’s 38%. With ZPD instructions, GPT-4o’s success rate increases to 45%, outperforming regular chain-of-thought prompting’s 34%. Our analysis systematically examines both syntactic and semantic errors, providing insights into the strengths and weaknesses of language models in generating domain-specific programs. [^1] Our resources can be found at https://github.com/zharry29/proc2pddl.
@@
Tianyi Zhang1thanks: Equal contribution.  Li Zhang1footnotemark:  Zhaoyi Hou3 Ziyu Wang1Yuling Gu2Peter Clark2 Chris Callison-Burch1Niket Tandon2 1University of Pennsylvania  2Allen Institute for Artificial Intelligence 3University of Pittsburg {zty|zharry|ccb}@upenn.edu
@@ 1 Introduction
Planning is the task of finding a sequence of actions to achieve a goal in a given environment Fikes and Nilsson (1971); LaValle (2006). In real life, the environment is often described with natural language texts. To enable text-based, automated planning, recent work has used language models (LMs) to generate plans Valmeekam et al. (2023a); Stein and Koller (2023). However, this approach is found to fall short with regard to both performance and interpretability Valmeekam et al. (2023c, b). Alternatively, another recent line of worked has instead used LMs to translate the natural language description of environments to planning domain definition language (PDDL) Ghallab et al. (1998). This symbolic representation can then be solved by a planner in a plan Collins et al. (2022); Lyu et al. (2023); Liu et al. (2023); Xie et al. (2023); Wong et al. (2023). Despite of the success of such a neurosymbolic method, all the above work has only been evaluated in closed-domains simulated environments such as a household (e.g., ALFRED Shridhar et al. (2020)) or discrete object placement (e.g., BlocksWorld Valmeekam et al. (2024)) (as shown in Table 1).
@@
Figure 2: Our formulation of the $\mathbb{DF}$ action prediction task is as follows: given a natural language procedure text and a domain file header, a language model (LM) follows Zone of Proximal Development (ZPD) instructions in three sequential skills to predict domain actions, including parameters, preconditions, and effects. During evaluation, the predicted $\mathbb{DF}$ is compared to a gold reference and used to solve corresponding $\mathbb{PF}$s.
To enable open-domain, text-based planning, we propose Proc2Pddl, a dataset to evaluate models’ ability to generate PDDL given procedural texts. Proc2Pddl consists of 27 pairs of open-domain procedures and PDDL representations. Each PDDL representation include a domain file $\mathbb{DF}$ that models the types, predicates, and actions, and a problem file $\mathbb{PF}$ that models the entities, initial states, and goal states, as illustrated in Figure 1. Because Proc2Pddl is not bound to any simulation, the PDDL representations are manually annotated by experts trained on this task to ensure validity, resulting in 27 domain files and 95 problem files.
Using this dataset, we study the task of action modeling Lindsay et al. (2017) formulated as follows. The input is some relevant natural language texts and the header of a $\mathbb{DF}$ (i.e., types, predicates, and names of actions). Based on a ZPD instruction, the output is the domain actions in the $\mathbb{DF}$ (i.e., parameters, preconditions, and effects). During evaluation, the predicted $\mathbb{DF}$ is 1) compared to a ground-truth $\mathbb{DF}$ as intrinsic evaluation, and 2) provided to a PDDL solver with ground-truth $\mathbb{PF}$s for the existence and correctness of plans as extrinsic evaluation. Our system is delineated in Figure 2. In this formulation, our assumption of the $\mathbb{DF}$ header is necessary to ensure the consistency of semantics between the $\mathbb{DF}$ and the $\mathbb{PF}$ for evaluation. It is also empirically motivated; for example, a kitchen robot may have access to the types like ‘ingredients’ and predicates like ‘diced’ via some information extraction system given descriptive texts, but it may still need to predict, for “swinging a knife”, the precondition that it is only safe to do so to the ‘ingredients’ and the effect that they will become ‘diced’.
Through our experiment, we show that the task of action modeling in Proc2Pddl is highly challenging to state-of-the-art LMs, where GPT-3.5 almost fails completely, GPT-4 can only generate exactly matching $\mathbb{DF}$s 16% of the time and solvable $\mathbb{PF}$s 33% of the time, and GPT-4o demonstrate 18% $\mathbb{DF}$s accuracy and 37% $\mathbb{PF}$s solving rate. By devising a ZPD instruction that prompt LMs to modularly generate PDDL through extraction-inference-translation approach, we improve action accuracy by 3% and problem solving by 2-7% . In our analysis, the syntactic errors indicate LMs’ weakness in generating low-resource and domain-specific programming languages Cassano et al. (2023) like PDDL, while the semantic errors suggest LMs’ inaccuracies to reason about actions and environments.
@@
2 Task Formulation
The task of predicting a planning domain definition in a text-based environment can be seen as translating natural language texts to PDDL symbolic language, which consists of a domain file ($\mathbb{DF}$) and one or more problem files ($\mathbb{PF}$s).
A $\mathbb{DF}$ defines all actions in the environment:
@@
A $\mathbb{PF}$ defines the initial and goal environments:
@@
Traditionally, the task of text-based PDDL generation involves predicting $\mathbb{PF}$ based on text $\mathbb{T}$, where a successfully generated $\mathbb{PF}$ can be solved by the predefined $\mathbb{DF}$.
In this paper, we address an alternative formulation, action modeling ($A$), in which the generated $\mathbb{DF}$, given text $\mathbb{T}$ and the domain header $H$ [^2] The domain header includes types, predicates, and names of actions in $\mathbb{DF}$. As the information specified by $H$ is guaranteed to be consistent with that of the $\mathbb{PF}$s, the evaluation is well-defined., is capable of producing plans for $\mathbb{PF}$s.
3 Dataset
We introduce the Proc2Pddl dataset of 27 different $\mathbb{T}$-$\mathbb{DF}$-$\mathbb{PF}$s tuples, drawing procedural texts from wikiHow articles of various topics (see Appendix A). A class of graduate students in a U.S. university with prior knowledge of PDDL are each given a wikiHow article and annotate a $\mathbb{DF}$ and multiple corresponding $\mathbb{PF}$s from the article, each with a gold plan to solve it. On average, there are 13.33 defined actions in a $\mathbb{DF}$ and 8.07 instantiated actions in a gold plan. In this work, all our data is used for evaluation, as all our methods are without task specific model training. Some sample data of Proc2Pddl can be found in Appendix B.
4 Methodology
We first introduce a novel prompt design option, ZPD, and then discuss the choices of text format ($\mathbb{T}$), which can range from 10 to 2,000 tokens and influence the selection of LMs.
4.1 ZPD Prompt Design
To predict domain actions $A$ based on relevant $\mathbb{T}$ and the header $H$, we prompt an LM in zero-shot or few-shot instructions. Our instruction employs Zone of Proximal Development (ZPD) theory proposed for human learning Vygotsky and Cole (1978), which is a variant of the chain-of-thought (CoT) approach. In typical CoT, a task is decomposed into several constituents (steps), i.e., parameters, precondition, and effect. In contrast, according to ZPD, the complex Proc2Pddl task is decomposed into atomic skills: 1) extracting the relevant description of an action; 2) extracting and inferring the incorporated entities and their state changes; and 3) translating the entity-state changes to accessible PDDL predicates. Next, we establish the relationships between these atomic skills: to perform the task, each skill is a prerequisite for the next. Finally, we explicitly instruct the LMs to incrementally perform the three basic skills, leading to the successful completion of the Proc2Pddl task (the prompt can be found in Appendix D):
  1. 1.
    Extraction: describe each action, including the expected preconditions and effects;
  2. 2.
    Inference: list the involved entities and their state changes;
  3. 3.
@@ 4.1 ZPD Prompt Design
Model % action acc. $\mathbb{PF}$ solve
gpt-3.5 0.2 1.0
gpt-4 15.9 33.7
+ CoT 9.3 21.1
+ ZPD 18.1 35.8
+ ZPD, 3 shot 11.9 23.2
gpt-4o 18.2 37.9
+ CoT 19.5 33.7
+ ZPD 21.4 45.3
+ ZPD, 3 shot 20.3 40.0
gold 100 100
Table 2: The intrinsic and extrinsic evaluation results for all main models. gpt-4(o) demonstrates non-trivial performance. With a ZPD instruction, the performance improves consistently.
@@
+ CoT 29.7 25 54.7
+ ZPD 42.2 29.7 48.1
gpt-4o 45.1 31.1 62.5
+ CoT 52.4 34.2 54.1
+ ZPD 53.5 40.1 53.5
Table 3: The generation accuracy of each component in actions has been evaluated. The ZPD instruction clearly aids in identifying implicit parameters (entities). Predicting preconditions is more challenging than predicting effects, as it requires a greater depth of implicit knowledge of entity states.
4.2 Choice of Input Text
We also consider the following choices of wikiHow text as $\mathbb{T}$.
($\mathbb{T}$ = all): All steps in a wikiHow article.
($\mathbb{T}$ = rel): In Proc2Pddl, each wikiHow article consists of step paragraphs that may or may not be used in defining the actions in the $\mathbb{DF}$. Hence, a mapping between actions and steps is also annotated. This context includes relevant steps to all actions in a $\mathbb{DF}$. (e.g., Step 1. Find fresh water… Step 2. Collect food… Step 7. Set up camp…)
($\mathbb{T}$ = map): Each action is mapped with steps based on the annotated mapping in Proc2Pddl.
(e.g., clean_water: Step 1. Find fresh water…)
($\mathbb{T}$ = sum): An one-line summary of each action annotated in Proc2Pddl.
(e.g., clean_water; boil water to clean it)
The four prompts are increasingly general. Distinguishing from the required skills, the full text condition demands accurate information extraction, while the text summary clearly defines the action but requires the model’s robust ability to infer implicit entity states. All prompts request an exact translation.
Intrinsic Extrinsic
Model % action acc. $\mathbb{PF}$ solve
w/o $\mathbb{T}$ (baseline) 13.7 26.3
$\mathbb{T}$ =sum 15.9 33.7
$\mathbb{T}$ =sum, ZPD 18.1 35.8
$\mathbb{T}$ =map 11.8 13.7
$\mathbb{T}$ =map, ZPD 8.9 26.3
$\mathbb{T}$ =rel 11.6 27.4
$\mathbb{T}$ =rel, ZPD 12.2 21.1
$\mathbb{T}$ =all 12.1 28.4
$\mathbb{T}$ =all, ZPD 12.1 31.6
Table 4: Performance of GPT-4 using different portions of text $\mathbb{T}$. Metrics include action-wide accuracy and the proportion of $\mathbb{PF}$s that can be solved.
4.3 Experiments
We conducted experiments with three large language models [^3] Due to the need for very long input and output, the choice of open-source models is limited. We are in progress of implementing Mixtral-8x7B.: GPT-3.5-turbo-16k, GPT-4-32k (dated June 2023), and GPT-4o. For GPT-4-32k, we used a maximum token limit of 10,000. GPT-3.5-turbo-16k and GPT-4o were tested with theirs default hyperparameters. The few-shot examples can be found in Appendix C.
@@ 5 Evaluation and Analysis
Now that a model generates the parameters, preconditions, and effects for each action, we have a complete $\mathbb{DF}$. We evaluate it in two ways (Figure 2). Intrinsically, we semantically compare the predicted $A$ with the ground-truth provided by our Proc2Pddl and report an action-wide accuracy. Equivalence of two action definitions does not depend on the naming of variables nor on the order within conjunctions (detailed in Appendix E). Extrinsically, to measure actions’ coherence, a BFS-based PDDL solver [^4] https://github.com/pucrs-automated-planning/pddl-parser attempts to solve ground-truth $\mathbb{PF}$s with the predicted $\mathbb{DF}$ and a success rate is reported. An unsolved $\mathbb{PF}$ is caused by (1.) no plan can be found, or (2.) the solver runs for more than 30 seconds, or (3.) the solver returns an error (usually a syntax error in the generated PDDL).
The intrinsic and extrinsic results are shown in Table 2. gpt-3.5-turbo which achieves impressive performance on many tasks has a close-to-zero performance. In contrast, gpt-4 performs significantly better with 18% action prediction accuracy and 36% solve rate of $\mathbb{PF}$s. The most advanced gpt-4o presents the highest performance, with 21% action accuracy and 45% $\mathbb{PF}$s solving rate. Still, the performance is far worse than ideal, showing that even a simplified open-domain planning formulation is challenging to state-of-the-art LMs.
ZPD Instruction Analysis
ZPD is helpful in each setting since it explicitly spells out many implicit entities and state changes in the inference stage which are critical to predicting parameters. In most situations, the model summarizes the action and extracts the entity states correctly, though sometimes missing a few implicit entities. However, ZPD’s bottleneck lies in the translation stage, during which there are mainly three types of errors.
@@
To address these, we leave to future work to demonstrate and standardize the translation process by clearly describing all necessary entity-state change and encouraging the model to compare and strictly match the given predicates. Finer-grained evaluation results are shown in Table 3 to tease out the performance regarding such component within an action. It is clear that the LM is worse at predicting preconditions than at predicting effects. This is understandable as procedural texts like wikiHow tend to be less explicit about predictions than about effects (e.g., from ‘bake for 10 minutes’ it is obvious that the food will be baked, but it is unclear what state it had been in).
Text Format Analysis
As shown in Table 4, in w/o $\mathbb{T}$ setting, fully relying on its implicit knowledge, the model is already capable of inferring PDDL syntactically and semantically. In w/ $\mathbb{T}$ settings, our model shows an ‘U’ performance in terms of the text length. Using a sentence-long description for each action ($\mathbb{T}$ = sum) provided by Proc2Pddl, the model achieves the best performance among all, showing a strong deduction ability with the limited but precise NL input. The $\mathbb{T}$ = all setting ensues, which requires the most extraction rather than inference. In contrast, the middle ones ($\mathbb{T}$ = rel/map) with decreasing signal-to-noise ratio lead to worse results, indicating its shortage of extraction-inference trade-off. The signals contain both the described entity states and step relations, explicitly and implicitly. This shortage may come less from the entity states (e.g., fish, spear in hunt_fish), but more from the relation between actions (e.g., make_spear to hunt_fish) which may be expressed in the $\mathbb{T}$ = sum and all settings.